# Models of interacting bosons with exact ground states: a unified approach

### v3

@@ Abstract
We define an infinite class of “frustration-free” interacting lattice quantum Hamiltonians for bosons, constructed such that their exact ground states have a density distribution specified by the Boltzmann weight of a corresponding classical lattice gas problem. By appropriately choosing the classical weights, we obtain boson representations of various known solvable models, including quantum dimer and vertex models, toric code, and certain Levin-Wen string-net models. We also systematically construct solvable models with other interesting ground states, including “quantum spin liquids,” supersolids, “Bose-Einstein insulators,” Bose liquids with “Bose surfaces”, and Bose-Einstein condensates that permit adiabatic evolution from a non-interacting limit to a Gutzwiller-projected limit.
@@
The models: The models we consider can be defined on any lattice in any dimension. In the present work, we will mainly consider systems that conserve boson number, but the scheme can be readily generalized to systems of bosons (e.g., magnons, phonons, or photons) whose number is not conserved, as discussed in App. E. The models are defined by Hamiltonians of the form:
@@
where $\delta K_{c}$ is the change of $K$ produced by the change in the bosonic density $\{n\}\to\{n\}_{c}$ produced by the operation of $B^{\dagger}_{c}$, i.e.
@@
where $\{n\}_{c}$ is related to $\{n\}$ by $n_{k}\to n_{k}+1\ \forall\ k\in c$.
To make the structure of the Hamiltonian clear, it is convenient to define the ‘soft projection’ operators (which are hermitian and non-negative)
@@
Ground states: The ground states of the Hamiltonian are a set of modified coherent states
@@
$\left(\hat{B}_{c}-\hat{B}_{c^{\prime}}\right)\hat{P}^{-1}|\alpha\rangle=\frac{1}{\mathcal{N}_{\alpha}}(\alpha^{L}-\alpha^{L})\exp\left[\alpha\sum_{i}\hat{b}^{\dagger}_{i}\right]\left|0\right\rangle,$ (10)
since the unmodified coherent states are eigenstates of the boson annihilation operators. It thus follows that $|\alpha\rangle$, which is a modified BEC state, is a ground state of $\hat{H}$, since each $\hat{H}_{c,c^{\prime}}$ is non-negative and $\hat{H}_{c,c^{\prime}}|\alpha\rangle=0$.
Since the system conserves boson number, one can also construct a ground state within each number sector. We discuss this construction and its properties in App. B.
@@
Clarifications: If the model is defined on a periodic lattice, the Hamiltonian so defined is translationally invariant provided both $\tau_{c,c^{\prime}}$ and $K$ are translationally invariant - but the construction works whether or not this symmetry is enforced. In particular, the ground-state is independent of $\tau_{c,c^{\prime}}$, which means that the ground-state can be translationally invariant even if the Hamiltonian is not. Conversely, if the classical problem defined by $K$ exhibits spontaneous translation symmetry breaking (i.e. some form of charge-density-wave - CDW - order), the corresponding ground-state will break the same symmetry, even if $\hat{H}$ is translationally invariant.
If both $\tau$ and $K$ are short-ranged, then, of course, so is $\hat{H}$. Under these conditions, barring the taking of a singular limit in which some parameter in $K$ tends to infinity (i.e. enforces a constraint), one generally expects that the connected correlators of the classical problem decay with distance, meaning that
@@
Note also that there is no need, anywhere in the above construction, to stick with clusters with a given character or number of operators, $L$. It is perfectly possible for $\hat{H}$ to simultaneously contain different types of terms hopping different numbers of bosons.
Dynamical density fluctuations: Besides the ground states, the dynamics of the density fluctuations can also be inferred. For concreteness, in this and the next section, we will consider the case of inversion symmetric systems on simple periodic lattices so that a single Bloch wave vector, $\bm{q}$, can label the density mode fluctuation.
The density operator at momentum $\bm{q}$,
@@
In the limit $\bm{q}\rightarrow\bm{0}$, for generic forms of short-range $\tau$ and $K$, it follows that $f(\bm{q})\sim\bm{q}^{2}$ and $s(\bm{q})\rightarrow s_{0}$ where $s_{0}\equiv\frac{\partial\bar{n}}{\partial\mu}$ is the compressibility of both the quantum and classical systems. Thus, so long as the classical system is compressible, the density excitation spectrum of the quantum system at $\bm{q}\rightarrow\bm{0}$ must vanish at least as fast as $\bm{q}^{2}$  [^2] This is consistent with several proven bounds for frustration-free systems [99, 100].
@@
For compressible systems, $s_{0}$ is finite while $f(\bm{q}\rightarrow 0)\sim\kappa\bm{q}^{2}$, so the long-wavelength limit of the theory is
@@
There are two interesting implications of this: 1) The first is that it provides an example of a system of strongly interacting bosons with ground-state properties that are still - in effect - equivalent to those of a non-interacting BEC. As already mentioned, the way to interpret this is as plotting a fine-tuned trajectory through Hamiltonian space lying precisely on the boundary of the superfluid phase - where almost any form of additional attractive interactions will trigger an instability to phase separation. 2) In the limit $u\rightarrow\infty$, the corresponding $\hat{P}$ operator becomes a Gutzwiller projection that constrains $n_{i}$ on each site to be either $p$ or $p+1$, where $p=\lfloor\bar{n}\rfloor$ is the largest integer that is no larger than the average particle number per site $\bar{n}$. This establishes the possibility of adiabatic evolution from the weakly interacting limit to a hard-core limit of the BEC ground state, despite arguments that the truncation of the Hilbert space (i.e. a change in the “exclusion statistics” [82]) that characterizes this limit must fundamentally change the emergent properties. An obvious exception in which the $u\to\infty$ limit does indeed change the character of the state arises in the case that $\bar{n}$ is an integer, in which case the classical system (and hence the quantum system) becomes incompressible in this limit, corresponding to a Bose-Mott insulating phase.  [^3] For $\bar{n}$ an integer and $u$ large but non-infinite, the system still possesses ODLRO, but with $\langle b_{k}\rangle\sim e^{-u}$ i.e. it is a form of “gossamer superfluidity.” [101]
@@
Other models with constrained Hilbert spaces: Further extending this methodology, one can also obtain ad hoc boson representations of certain Levin-Wen string-net models [8, 9], Kitaev’s toric code [7] or quantum vertex models [86, 13, 87, 79]. The strategy is to introduce several flavors of bosons on the lattice edge to represent the possible ‘string types’ and then use an appropriate $K(\{n\})$ to impose a single-occupancy constraint on each edge and a constraint at each vertex so that surrounding string types obey the ‘branching rules.’ By considering the quantum resonance of the strings consistent with the rules, one can, in this way, naturally obtain the solvable Hamiltonians describing certain deconfined gauge theories.
@@
Many aspects of systems with this sort of kinetic energy can be treated from the present perspective by suitably defining the clusters $c$ and $c^{\prime}$ connected by $\hat{H}_{c,c^{\prime}}$. The resulting solvable ground states can still be compressible and possess ODLRO. However, for processes that conserve the center-of-mass position of the bosons, $g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})$ vanishes at least as $\bm{q}^{2}$ as $\bm{q}\to\bm{0}$, implying that $f(\bm{q})\sim|\bm{q}|^{4}$ as $\bm{q}\to\bm{0}$ (see Eq. 18). It follows from Eq. 20 that the superfluid stiffness (i.e. the coefficient of $\left({\bf\nabla}\theta\right)^{2}$ in the effective action) vanishes. Since any uniform current-carrying state would violate the assumed dipole conservation, such states are also, in some sense, insulating - they have thus been named ‘Bose-Einstein insulators [93]’.
Bose liquids with ‘Bose surfaces’: When the non-zero $\tau_{c,c^{\prime}}$’s all correspond to $\left|g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})\right|^{2}$ that vanish on a submanifold $\mathcal{M}$ of the momentum space, this implies that $f(\bm{q})$ as well as $E_{\text{SMA}}(\bm{q})$ (assuming the system is compressible) vanish on $\mathcal{M}$. For most cases, $\mathcal{M}$ is a single point at $\bm{q}=\bm{0}$, but sometimes it can be higher dimensional. As a concrete example, consider short-range resonant pair hopping on a square lattice where a pair of bosons from one set of opposite vertices, $c$, of an elementary plaquette is shifted to the complementary vertices, $c^{\prime}$. [Note that this process happens to be dipole conserving; moreover, if we identify the bosons with dimers and the lattice vertices with the edge centers of another rotated square lattice, this is the sort of dimer resonance term mentioned above that must be introduced to obtain the dynamical terms in the QDM at $\bar{N}=1$.] The resulting $f(\bm{q})\sim\sin^{2}q_{x}\sin^{2}q_{y}$ not only vanishes as $\bm{q}\rightarrow\bm{0}$, but also along lines in $\bm{q}$-space. Therefore, the system has gapless modes along one-dimensional ‘Bose-surfaces’ [94] in the Brillouin zone.
@@
Discussion: In this work, we have developed a unified approach for constructing “frustration-free” Hamiltonians for number-conserving bosons whose exact ground-state properties can be inferred from the properties of an associated classical statistical mechanical problem on the same lattice. For the known solvable models, our scheme provides not only an alternative representation that might be realizable on engineered platforms like optical lattices but also a natural route from a strongly constrained problem to a free (trivial) limit of the boson problem. When the constrained limit of the problem exhibits an exotic quantum phase of matter, this evolution must drive exotic transitions out of them, whose properties may be studied with these engineered models. The critical points constructed this way are “conformal quantum critical points” [13] with unusual dynamical properties [16, 96].
The asymptotic behaviors of correlations in $d$ dimensional quantum ($T=0$) and classical ($T>0$) systems are generically different. Thus, solvable models of the sort we have discussed are always, to some degree, non-generic. Indeed, in many cases, as we have shown, the solvable models describe systems on the phase boundary beyond which there is instability - often to phase separation. A dramatic manifestation of this point is that the charge $U(1)$ symmetry is broken in the ground state of such models even in one dimension, similar to certain frustration-free spin models [97]. Still, it is possible to infer generic properties of interesting phases by constructing an effective field theoretical representation of the solvable model and then addressing the effect of symmetry-allowed perturbations.
There are a variety of further directions that presumably can be successfully explored by extensions of the present approach. For one thing, disordered versions of the present models can easily be constructed: An artificial version of this (similar to the Mattis model [98]) is achieved by choosing random hopping matrix elements, $\tau_{c,c^{\prime}}$, while keeping $K$ translationally invariant, which results in a translationally invariant ground state, but where the disorder is strongly reflected in the excitation spectrum. More interestingly, if $K$ is not translationally invariant, then the resulting potential energy will not be either, nor will the associated ground state, despite still exhibiting ODLRO.
@@ Appendix A Generalized Witten’s conjugation method
The specific class of models discussed in the main text is of this form with $\hat{\mathcal{O}}_{l}\rightarrow\hat{\mathcal{B}}_{c,c^{\prime}}$, $\hat{\mathcal{P}}\rightarrow\hat{P}$, and ^$\hat{\mathcal{M}}_{l}\rightarrow\hat{P}_{c}\hat{P}_{c^{\prime}}$, the later operator so chosen as to ensure that the resulting hopping terms retains a simple and physical form. It is also similar to a construction for interacting boson models in the continuum studied in Ref. [14] where the engineered Hamiltonian is written as
$\hat{\mathcal{H}}=$ $\hat{P}^{-1}\left(\nabla\psi^{\dagger}\right)\hat{P}\cdot\hat{P}\left(\nabla\psi\right)\hat{P}^{-1}.$ (28)
@@ Appendix B The number basis
$\mathcal{N}_{M}^{2}=$ $\left\langle 0\right|\left(\sum_{i}\hat{b}_{i}\right)^{M}\hat{P}^{2}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}\left|0\right\rangle$ (31)
$=$ $\sum_{\{n\}|\sum_{k}n_{k}=M}{M\choose n_{1},\dots,n_{\mathsf{V}}}^{2}\mathrm{e}^{-K(\{n\})}$
@@ Appendix C The mapping to classical problem
$\left|\alpha\right\rangle=$ $\frac{1}{\mathcal{N}_{\alpha}}\hat{P}\sum_{M=0}^{\infty}\exp\left[\alpha\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)\right]\left|0\right\rangle$ (35)
$=$ $\frac{1}{\mathcal{N}_{\alpha}}\hat{P}\sum_{M=0}^{\infty}\frac{\alpha^{M}}{M!}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}\left|0\right\rangle$ (36)
@@
$=$ $\frac{\delta_{L,L^{\prime}}}{\mathcal{N}_{M}}\sum_{\{n\}|\sum_{k}n_{k}=M-L}\frac{1}{\prod_{k}(n_{k}!)}\langle 0|\left(\sum_{i}\hat{b}_{i}\right)^{M}\hat{P}\hat{B}^{\dagger}_{c}\prod_{k}(\hat{b}^{\dagger}_{k})^{n_{k}}|0\rangle\langle 0|\prod_{k}(\hat{b}_{k})^{n_{k}}\hat{B}_{c^{\prime}}\hat{P}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}|0\rangle$ (43)
$=$ $\frac{\delta_{L,L^{\prime}}}{\mathcal{N}_{M}}\sum_{\{n\}|\sum_{k}n_{k}=M-L}\frac{(M!)^{2}}{\prod_{k}(n_{k}!)}\mathrm{e}^{-\delta K_{c}(\{n\})/2-\delta K_{c^{\prime}}(\{n\})/2}\mathrm{e}^{-K(\{n\})}$ (44)
$=$ $\frac{\delta_{L,L^{\prime}}\mathcal{Z}_{M-L}}{\mathcal{Z}_{M}}\Big{\langle}\!\!\Big{\langle}\mathrm{e}^{-\delta K_{i_{1}\dots i_{L}}(\{n\})/2}\mathrm{e}^{-\delta K_{j_{1}\dots j_{L}}(\{n\})/2}\Big{\rangle}\!\!\Big{\rangle}_{M-L}$ (45)

### v4

@@ Abstract
We define an infinite class of “frustration-free” interacting lattice quantum Hamiltonians for bosons, constructed such that their exact ground states have a density distribution specified by the Boltzmann weight of a corresponding classical lattice gas problem. By appropriately choosing the classical weights, we obtain boson representations of various known solvable models, including quantum dimer and vertex models, toric code, and certain Levin-Wen string-net models. We also systematically construct solvable models with other interesting ground states, including quantum spin liquids, supersolids, “Bose-Einstein insulators,” Bose liquids with “Bose surfaces”, and Bose-Einstein condensates that permit adiabatic evolution from a non-interacting limit to a Gutzwiller-projected limit.
@@
The models: The models we consider can be defined on any lattice in any dimension. In the present work, we will consider systems that conserve boson number, but the scheme can be readily generalized to systems of bosons whose number is not conserved (e.g., spin waves, phonons, or photons), as discussed in App. E. The models are defined by Hamiltonians of the form:
@@
where $\delta K_{c}$ is the change of $K$ produced by the change in the configuration corresponding to the operation of $B^{\dagger}_{c}$,
@@
where $\{n\}_{c}$ is defined as the occupation configuration of $\hat{B}_{c}^{\dagger}|\{n\}\rangle$.
To make the structure of the Hamiltonian clear, it is convenient to define a set of ‘soft projection’ operators (which are non-negative, and mutually commuting)
@@
Note also that there is no need, anywhere in the above construction, to stick with clusters with a given character or number of operators, $L$. It is perfectly possible for $\hat{H}$ to simultaneously contain different types of terms hopping different numbers of bosons. Moreover, if both $\tau$ and $K$ are short-ranged, then, of course, so is $\hat{H}$.
Ground states: The ground states of the Hamiltonian are a set of coherent states modified by $\hat{P}$
@@
$\left(\hat{B}_{c}-\hat{B}_{c^{\prime}}\right)\hat{P}^{-1}|\alpha\rangle$
$=$ $\frac{1}{\mathcal{N}_{\alpha}}(\alpha^{L}-\alpha^{L})\exp\left[\alpha\sum_{i}\hat{b}^{\dagger}_{i}\right]\left|0\right\rangle=0.$ (10)
It thus follows that $|\alpha\rangle$, which is a modified BEC state, is a ground state of $\hat{H}$, since each $\hat{H}_{c,c^{\prime}}$ is non-negative and $\hat{H}_{c,c^{\prime}}|\alpha\rangle=0$. Since the system conserves boson number, one can also construct a ground state within each number sector. We discuss this construction and its properties in App. B.
@@
It is interesting to note that the ground-states are determined by $K(\{n\})$ and independent of $\tau_{c,c^{\prime}}$, which means that the ground states can have higher symmetry than the Hamiltonian. As a dramatic example, one can construct a disordered Hamiltonian with spatially random hopping matrix elements, $\tau_{c,c^{\prime}}$, while keeping $K(\{n\})$ translationally invariant, which results in a uniform ground state (similar to the Mattis model [72]).
@@
When $K$ is short-ranged, barring the taking of a singular limit in which some parameter in $K$ tends to infinity (i.e. enforces a constraint), one generally expects that the connected correlators of the classical problem decay with distance, meaning that
@@
Dynamical density fluctuations: Besides the ground states, the dynamics of the density fluctuations can also be inferred. For concreteness, in this and the next section, we will consider the case of inversion symmetric systems on simple periodic lattices so that a single Bloch wave vector, $\bm{q}$, can label the density mode fluctuation. The density operator at momentum $\bm{q}$,
@@
In the limit $\bm{q}\rightarrow\bm{0}$, for generic forms of short-range $\tau$ and $K$, it follows that $f(\bm{q})\sim\bm{q}^{2}$ and $s(\bm{q})\rightarrow s_{0}$ where $s_{0}\equiv\frac{\partial\bar{n}}{\partial\mu}$ is the compressibility of the classical systems. Thus, so long as the classical system is compressible, the density excitation spectrum of the quantum system at $\bm{q}\rightarrow\bm{0}$ must vanish at least as fast as $\bm{q}^{2}$  [^2] This is consistent with several proven bounds for frustration-free systems [100, 101].
@@
For compressible classical systems, $s_{0}$ is finite while $f(\bm{q}\rightarrow 0)\sim\kappa\bm{q}^{2}$, so the long-wavelength limit of the theory is
@@
There are two interesting implications of this: 1) The first is that it provides an example of a system of strongly interacting bosons with ground-state properties that are still - in effect - equivalent to those of a non-interacting BEC. As already mentioned, the way to interpret this is as plotting a fine-tuned trajectory through Hamiltonian space lying precisely on the boundary of the superfluid phase - where almost any form of additional attractive interactions will trigger an instability to phase separation. 2) In the limit $u\rightarrow\infty$, the corresponding $\hat{P}$ operator becomes a Gutzwiller projection that constrains $n_{i}$ on each site to be either $p$ or $p+1$, where $p=\lfloor\bar{n}\rfloor$ is the largest integer that is no larger than the average particle number per site $\bar{n}$. This establishes the possibility of adiabatic evolution from the weakly interacting limit to a hard-core limit of the BEC ground state, despite arguments that the truncation of the Hilbert space (i.e. a change in the “exclusion statistics” [83]) that characterizes this limit must fundamentally change the emergent properties. An obvious exception in which the $u\to\infty$ limit does indeed change the character of the state arises in the case that $\bar{n}$ is an integer, in which case the classical system becomes incompressible in this limit, corresponding to a Bose-Mott insulating phase.  [^3] For $\bar{n}$ an integer and $u$ large but non-infinite, the system still possesses ODLRO, but with $\langle b_{k}\rangle\sim e^{-u}$ i.e. it is a form of “gossamer superfluidity.” [102]
@@
Other models with constrained Hilbert spaces: Further extending this methodology, one can also obtain boson representations of certain Levin-Wen string-net models [8, 9], Kitaev’s toric code [7] or quantum vertex models [87, 13, 88, 80]. The strategy is to introduce several flavors of bosons on the lattice edge to represent the possible ‘string types’ and then use an appropriate $K(\{n\})$ to impose a single-occupancy constraint on each edge and a constraint at each vertex so that surrounding string types obey the ‘branching rules.’ By considering the quantum resonance of the strings consistent with the rules, one can, in this way, naturally obtain certain solvable Hamiltonians for deconfined gauge theories, whose groundstates have trivial relative phases between different string configurations.
@@
Many aspects of systems with this sort of kinetic energy can be treated from the present perspective by suitably defining the clusters $c$ and $c^{\prime}$ connected by $\hat{H}_{c,c^{\prime}}$. The resulting solvable ground states can still possess ODLRO. However, for processes that conserve the center-of-mass position of the bosons, $g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})$ vanishes at least as $\bm{q}^{2}$ as $\bm{q}\to\bm{0}$, implying that $f(\bm{q})\sim|\bm{q}|^{4}$ as $\bm{q}\to\bm{0}$ (see Eq. 18). It follows from Eq. 20 that the superfluid stiffness (i.e. the coefficient of $\left({\bf\nabla}\theta\right)^{2}$ in the effective action) vanishes. Since any uniform current-carrying state would violate the assumed dipole conservation, such states are also, in some sense, insulating - they have thus been named ‘Bose-Einstein insulators [94]’.
Bose liquids with ‘Bose surfaces’: When the non-zero $\tau_{c,c^{\prime}}$’s all correspond to $\left|g_{c}(\bm{q})-g_{c^{\prime}}(\bm{q})\right|^{2}$ that vanish on a submanifold $\mathcal{M}$ of the momentum space, this implies that $f(\bm{q})$ as well as $E_{\text{SMA}}(\bm{q})$ (assuming the classical system is compressible) vanish on $\mathcal{M}$. For most cases, $\mathcal{M}$ is a single point at $\bm{q}=\bm{0}$, but sometimes it can be higher dimensional. As a concrete example, consider short-range resonant pair hopping on a square lattice where a pair of bosons from one set of opposite vertices, $c$, of an elementary plaquette is shifted to the complementary vertices, $c^{\prime}$. [Note that this process happens to be dipole conserving; moreover, if we identify the bosons with dimers and the lattice vertices with the edge centers of another rotated square lattice, this is the sort of dimer resonance term mentioned above that must be introduced to obtain the dynamical terms in the QDM at $\bar{N}=1$.] The resulting $f(\bm{q})\sim\sin^{2}q_{x}\sin^{2}q_{y}$ not only vanishes as $\bm{q}\rightarrow\bm{0}$, but also along lines in $\bm{q}$-space. Therefore, the system has gapless modes along one-dimensional ‘Bose-surfaces’ [95] in the Brillouin zone.
@@
Discussion: For the known solvable models, our scheme provides not only an alternative representation that might be realizable on engineered platforms like optical lattices but also a natural route from a strongly constrained problem to a free (trivial) limit of the boson problem. When the constrained limit of the problem exhibits an exotic quantum phase of matter, this evolution must drive exotic transitions out of them, whose properties may be studied with these engineered models. The critical points constructed this way are “quantum Lifshitz points” [13] with unusual dynamical properties [16, 97].
The asymptotic behaviors of correlations in $d$ dimensional quantum ($T=0$) and classical ($T>0$) systems are generically different. Thus, solvable models of the sort we have discussed are always, to some degree, non-generic. Indeed, in many cases, as we have shown, the solvable models describe systems on the phase boundary beyond which there is an instability - often to phase separation. A dramatic manifestation of this point is that the charge $U(1)$ symmetry is broken in the ground state of such models even in one dimension, similar to certain frustration-free spin models [98] and “transverse quantum fluids” [99]. Still, it is possible to infer generic properties of interesting phases by constructing an effective field theoretical representation of the solvable model and then addressing the effect of symmetry-allowed perturbations.
A limitation of the specific construction in the current work is that the relative phase between different configurations is trivial, which can be overcome by considering complex-valued $K(\{n\})$ functions and/or starting from other frustration-free boson hopping models with zero mode(s) with non-trivial phase structures. We will leave exploration of these generalizations to future work.
@@ Appendix A Generalized Witten’s conjugation method
The specific class of models discussed in the main text is of this form with $\hat{\mathcal{O}}_{l}\rightarrow\hat{B}_{c}-\hat{B}_{c^{\prime}}$, $\hat{\mathcal{P}}\rightarrow\hat{P}$, and ^$\hat{\mathcal{M}}_{l}\rightarrow\hat{P}_{c}\hat{P}_{c^{\prime}}\hat{P}^{2}$, the later operator so chosen as to ensure that the resulting hopping terms retains a simple and physical form. It is also similar to a construction for interacting boson models in the continuum studied in Ref. [14] where the engineered Hamiltonian is written as
$\hat{\mathcal{H}}=$ $\hat{P}^{-1}\left(\nabla\hat{\psi}^{\dagger}\right)\hat{P}\cdot\hat{P}\left(\nabla\hat{\psi}\right)\hat{P}^{-1}.$ (28)
@@ Appendix B The number basis
$\mathcal{N}_{M}^{2}=$$\sum_{\{n\}|\sum_{k}n_{k}=M}{M\choose n_{1},\dots,n_{\mathsf{V}}}^{2}\mathrm{e}^{-K(\{n\})}$
@@ Appendix C The mapping to classical problem
$\left|\alpha\right\rangle=$ $\frac{1}{\mathcal{N}_{\alpha}}\hat{P}\sum_{M=0}^{\infty}\frac{\alpha^{M}}{M!}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}\left|0\right\rangle$ (34)
@@
$=$ $\frac{\delta_{L,L^{\prime}}}{\mathcal{N}^{2}_{M}}\sum_{\{n\}|\sum_{k}n_{k}=M-L}\frac{1}{\prod_{k}(n_{k}!)}\langle 0|\left(\sum_{i}\hat{b}_{i}\right)^{M}\hat{P}\hat{B}^{\dagger}_{c}\prod_{k}(\hat{b}^{\dagger}_{k})^{n_{k}}|0\rangle\langle 0|\prod_{k}(\hat{b}_{k})^{n_{k}}\hat{B}_{c^{\prime}}\hat{P}\left(\sum_{i}\hat{b}^{\dagger}_{i}\right)^{M}|0\rangle$ (44)
$=$ $\frac{\delta_{L,L^{\prime}}}{\mathcal{N}^{2}_{M}}\sum_{\{n\}|\sum_{k}n_{k}=M-L}\frac{(M!)^{2}}{\prod_{k}(n_{k}!)}\mathrm{e}^{-\delta K_{c}(\{n\})/2-\delta K_{c^{\prime}}(\{n\})/2}\mathrm{e}^{-K(\{n\})}$ (45)
$=$ $\frac{\delta_{L,L^{\prime}}\mathcal{Z}_{M-L}}{\mathcal{Z}_{M}}\Big{\langle}\!\!\Big{\langle}\mathrm{e}^{-\delta K_{c}(\{n\})/2}\mathrm{e}^{-\delta K_{c^{\prime}}(\{n\})/2}\Big{\rangle}\!\!\Big{\rangle}_{M-L}$ (46)